A 76 Netwon traffic light is supported at the midpoint of a 27 meter length of cable between two poles. Find the tension in each cable segment if the cable sags a vertical distance of 1 m.

Tension =  [Num]Newtons
Answer #1

h = 1 m

L = 27 m

l = L/2 = 27/2 = 13.5 m

2 T sin(theta) = 76 N

2 * T * 1 / 13.5 = 76

T = 76 * 13.5 /2 = 513 N

=======================================

Tension = 513 Newtons